# Electrically tunable High-Chern-number quasiflat bands in twisted antiferromagnetic topological insulators

### v1

@@ V Summary and discussion
In summary, based on the effective model analysis, we have proposed realizing high-Chern-number ($C=n$) flat band states near the Fermi level in twisted AFM TI thin films preserving an $n$-fold rotational symmetry. The intertwined Dirac-cone states induced by anisotropic inter-surface coupling are found to play a crucial role in forming these nontrivial flat bands through the twisting procedure. Furthermore, we have also demonstrated that an out-of-plane electric field could drive topological phase transitions between nontrivial ($C=n$) and trivial ($C=0$) flat band states. Our work not only sheds light on the significance of the recently proposed intertwined Dirac-cone states, but also open a new avenue to realize high-Chern-number flat bands.
@@ Acknowledgements.
This work is supported by National Key Projects for Research and Development of China (Grants No.2021YFA1400400, No.2017YFA0303203 and 2023YFA1407001), the Fundamental Research Funds for the Central Universities (Grant No. 020414380185), Natural Science Foundation of Jiangsu Province (No. BK20200007), the Natural Science Foundation of China (No. 12074181, No. 12104217, and No. 11834006), the Fok Ying-Tong Education Foundation of China (Grant No. 161006), and the Department of Science and Technology of Jiangsu Province (BK20220032)

### v2

@@ V Summary and discussion
In summary, based on the effective model analysis, we have proposed realizing high-Chern-number ($C=n$) flat-band states near the Fermi level in twisted AFM TI thin films preserving an $n$-fold rotational symmetry. The intertwined Dirac-cone states induced by anisotropic inter-surface coupling are found to play a crucial role in forming these nontrivial flat bands through the twisting procedure. Furthermore, we have also demonstrated that an out-of-plane electric field could drive topological phase transitions between nontrivial ($C=n$) and trivial ($C=0$) flat-band states. Our work not only sheds light on the significance of the recently proposed intertwined Dirac-cone states, but also open a new avenue to realize high-Chern-number flat bands.
@@ Acknowledgements.
This work is supported by National Key Projects for Research and Development of China (Grants No. 2021YFA1400400 and No. 2023YFA1407001), the Fundamental Research Funds for the Central Universities (Grant No. 020414380185), Natural Science Foundation of Jiangsu Province (No. BK20200007), the Natural Science Foundation of China (No. 12074181, No. 12104217, and No. 92365203), and the Department of Science and Technology of Jiangsu Province (BK20220032)